OpenStudy (anonymous):

explain the steps you take to write the ratio 5:25 in simplest form.

OpenStudy (mayankdevnani):

\[\huge \bf \frac{5}{25}\] you have to write in simplest form, see that any factor common between 5 and 25 ???

OpenStudy (mayankdevnani):

@zaqwsxedcrfv

OpenStudy (anonymous):

tell me the steps

OpenStudy (mayankdevnani):

tell me what is factor common between 5 and 25 ???

OpenStudy (anonymous):

to write in simplest form

OpenStudy (anonymous):

wait

OpenStudy (mayankdevnani):

take your time

OpenStudy (anonymous):

1 and 5

OpenStudy (mayankdevnani):

correct

OpenStudy (texaschic101):

ratios can be written as fractions...for instance, 2 : 3 can also be written as 2/3. So your ratio 5:25 can be written as 5/25 which can be reduced to 1/5...which in turn can be changed back to a ratio of 1 : 5 in simplest form.

OpenStudy (mayankdevnani):

but when we divide it by 1 ,we can get its simplest form. Right? @zaqwsxedcrfv

OpenStudy (anonymous):

so that the steps

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

the steps are basically reducing a fraction and turning it into a ratio

OpenStudy (anonymous):

oh i get ir

OpenStudy (mayankdevnani):

\[\large \bf \frac{5}{25}\] there is one common factor i.e 5 so when we divide it,the fraction converts into SIMPLEST FORM.

OpenStudy (texaschic101):

just reduce it

OpenStudy (anonymous):

k

OpenStudy (mayankdevnani):

so,divide numerator and denominator by 5, \[\huge \bf \frac{\frac{5}{5}}{\frac{25}{5}}=\frac{1}{5}=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

understood? @zaqwsxedcrfv

OpenStudy (anonymous):

yes

OpenStudy (mayankdevnani):

good :)

OpenStudy (anonymous):

want to be my friends

OpenStudy (anonymous):

i am your fan

OpenStudy (mayankdevnani):

why not

OpenStudy (anonymous):

add me